# VERUS-LM: a Versatile Framework for Combining LLMs with Symbolic Reasoning

### v1

@@ 5.1 Detect Reasoning Task
    Model Generation/Expansion: generate $n$ logical models of the theory $T$ (i.e., structures $S$ such that $S\models T$). Possibly, the interpretation $\sigma^{S}$ that some of the symbols $\sigma$ should have in this model $S$ is already given (e.g., we might provide the value for constants car_value and car_type and ask the engine to fill in the interpretation for the remaining symbols).
@@
    Satisfiability: verify if at least one such model $S$ exists for a given theory $T$.
@@
    Logical entailment: verify whether a given statement $\phi$ is logically entailed by the theory $T$.
@@
Detecting the correct inference to be executed is crucial to ensure that questions are correctly answered. To detect the intended inference from the a question, we fine-tuned a BERT sentence classifier Reimers and Gurevych (2019) on a custom dataset, which contains over 1000 natural language questions and their corresponding logical reasoning task.
@@ 5.2 Information Extraction
Typically, in addition to a theory $T$, an interpretation is given for at least some of the symbols of $T$. For instance, the problem statement may be: “Given that Ann is 16 and Barbara is 32, who is eligible for insurance?” Such information should be extracted and included as part of a structure on which the reasoning is performed (in this case, $age^{S}=\{\text{Ann}\mapsto 16,\text{Barbara}\mapsto 32\}$). In addition, some other reasoning tasks also require additional information of similar complexity, e.g., optimization requires that we know which term $t$ should be minimized / maximized.
This information extraction task is made significantly easier by the fact that, at this point, we already know the vocabulary. We can therefore use language model grammars, such as llama.cpp GNBF grammar Gerganov (2023), to force the output to adhere to an automatically generated KB-specific grammar, ensuring that only correctly typed interpretations are generated. As we will show, this pruning of the output space allows us to achieves the same performance for this task with a small language model (SLM) as with a larger LLM. Since such a SLM can run on a standard laptop, this may significantly reduce computational costs, as well as enhance data security, by avoiding the transfer of potentially sensitive information (e.g., the personal data of customers who apply for insurance) to external servers.
@@
We currently do not attempt to automatically detect whether a specific question contains complex formulas or not, and thus requires an LLM, beacuse in different applications, we may wish to make this decision in different ways. If we do not care about computational cost, we can just always invoke the LLM for optimal accuracy. Alternatively, we could base the decision whether to use an SLM or LLM purely on the detected reasoning task (e.g., Entailment is typically more likely to need an LLM), or we could even train a specific classifier to make the decision.
@@ 6.1 Validation
We first conduct a basic validation of VERUS-LM by checking (1) whether our use of a reasoning engine indeed leads to better perfomance on logical reasoning tasks than a stand-alone LLMs and (2) that the VERUS-LM pipeline is indeed able to correctly identify our eight different reasoning tasks from natural language text. Since existing datasets typically cover only one or at most a few reasoning tasks, we constructed our own dataset called DivLR in which all eight reasoning tasks are present. It consists of 115 questions covering six domains: investment Strategies, COVID restrictions, water Irrigation needs, a Handyman and two variants of the Body Mass Index (BMI) domain. The reason for having two BMI variants is that information about BMI (i.e., how to calculate it, the different risk levels associated to BMI values, etc.) probably occurs in LLM training data. In addition to our BMI domain which contains the well-known definition, we also defined a B* domain in which novel formulas and ranges for BMI were made up.
We compare VERUS-LM to baseline language models on this dataset. We compare with both an SLM and an LLM, and consider two versions of our pipeline: one in which our information extraction is done by an SLM (V-SLM) and one in which this is done by an LLM (V-SLM). the LLM is Gemini 1.5 Pro Team Gemini et al. (2023) and the SLM is a quantised version of Phi 3.5 Abdin et al. (2024), a 3.8-billion-parameter language model whose inference cost is about 10,000 times lower than large models like GPT-4. All experiments were performed on a Mac M2 Pro CPU with 32GB of RAM.
@@
Table 1: Result of VERUS-LM with SLM and LLM for information extraction and SLM and LLM separately as baseline. SLM = Phi 3.5 mini instruct, LLM = Gemini 1.5 Pro
Table 1 shows that both V-SLM and V-LLM consistently outperform both the SLM and the LLM on their own. However, V-SLM performs significantly worse on the two BMI domains than on the other domain, and also significantly worse that V-LLM on these domains. A deeper analysis shows that the SLM typically refuses to simply extract the information that the reasoning engine needs, and instead tries to perform BMI-calculations itself. We conjecture that this is due to the BMI-related examples that are likely part of the data on which the SLM was trained.
A similar but less striking issue can be seen when comparing the results of the LLM on both the B and B* domain: the observed drop in accuracy (66.7$\rightarrow$50) can be explained by the fact that the LLM has memorised the correct definition of BMI from its training data, and that is unable reason with a different definition, when one is explicitly provided. This suggests that, more generally, if an LLM at some point learns out-dated facts, it is hard to correct these in the prompt, without retraining the network. We see that same issue is still somewhat observable in V-LLM (88.9$\rightarrow$83.3), but here the reasoning engine pipeline substantially reduces the effect.
@@
Table 2 confirms that our BERT classifier identifies the correct logical reasoning task with an average accuracy of 92.6%, performing consistently well on all tasks (with a small dip for Propagation). For the subsequent information extraction, we see that, for most of the reasoning tasks, V-LLM and S-SLM perform about equally well. The biggest difference is seen for Entailement. This is the only reasoning task in our DivLR dataset for which complex formulas need to be constructed at inference time and V-SLM predictably struggles with this (accuracy 29.6%). For the Propagation task, we also observe a difference in accuracy (-25.6%) between V-SLM and V-LLM. This is primarily due V-SLM struggling with the BMI domain, as already discussed.
Overall, we conclude that the proposed pipeline makes sense: (1) introducing a logical reasoning engine into the pipeline indeed leads to better performance then using only a language model; (2) different logical reasoning tasks can indeed by successfully identified by a simple BERT classifier; (3) all of the necessary information for the reasoning tasks can indeed be extracted with a small language model, as long as no complex formulas need to be constructed, and with the caveat that an SLM might be less robust when handling concepts that already occurred frequently in its training data.
@@ 6.2 Comparison to state-of-the-art
Dataset Example At inference time # Options Reasoning Task #
PrOntoQA Saparov and He (2022) ”Stella is not hot.” Only Data 2: True,False Satisfiability 500
[…]
Table 3: Overview datasets used for comparison with state-of-the-art
System ProntoQA Proofwriter Folio Logical Deduction AR-LSAT Average ST Dev
@@
Table 4: Comparison of the results of the neurosymbolic system. Results for Logic-LM, SymboCOT and LINC come from their respective papers Pan et al. (2023); Xu et al. (2024); Olausson et al. (2023), and the LLM baselines originate from Xu et al. (2024) as it is the most recently published paper. The results for Logic-LM are those with self refinements included, but without the LLM fallback in case of recurring syntax error. The best results per dataset are put in bold, the second best results are italicized.
@@
We compare VERUS-LM against three LLM-based approaches and two neurosymbolic approaches for logical reasoning: (1) GPT-4 answering directly, (2) GPT-4 with Chain-of-Thought, (3) Linc Olausson et al. (2023) (only FOLIO and ProofWriter), (4) Logic-LM Pan et al. (2023) and (5) SymbCot Xu et al. (2024).
As shown in Table 4, all systems achieve a roughly similar performance on the first four benchmarks, with VERUS-LM scoring at most 5% worse than the best system on each. For the challenging AR-LSAT benchmark, VERUS-LM significantly outperforms all other methods, doing about 25% better than the second best system. We believe that this is due to (1) the generality of our approach, which does not use task-specific examples for in-context learning; (2) FO($\cdot$)’s expressiveness which allows for, e.g., straightforward representations of aggregates, such as “At least three CDs are on sale”: #$\#\{c\in\textit{CD}:on\_sale(c)\}\geq 3$; (3) our support for different forms or reasoning, four of which were detected by the classifier and executed, as shown in Table 5.
Reasoning task Distribution Exe_Acc
@@
Table 5: The distribution of reasoning tasks detected in AR-LSAT dataset
@@ 6.3 Effect of Self-Refinement
[…]
ProofWriter Syntax 92.3 $\uparrow$ 2.3 95.7 $=$ 88.3 $\uparrow$ 2.1
@@ 7 Limitations
Though VERUS-LM expands on the state of the art, it faces some limitations. Certain tasks are better suited to imperative programming paradigms rather than declarative approaches. Additionally, the expressiveness of FO($\cdot$) has limitations, making it among others unsuitable for representing higher-order logic. Due to VERUS-LM’s relience on a logical reasoning engine, it also inherits some of its limitations, such as SAT problems in very large domains. Furthermore, though the KB-creation of VERUS-LM is based on a generic prompting pipeline, its accuracy in entirely different logical paradigms remains a uncertain.

### v2

@@ 5.1 Detect Reasoning Task
    Model Generation/Expansion: generate $n$ logical models of the theory $T$ (i.e., structures $S$ such that $S\models T$). Possibly, the interpretation $\sigma^{S}$ that some of the symbols $\sigma$ should have in this model $S$ is already given (e.g., we might provide the value for constants car_value and car_type and ask the reasoner to complete the interpretation for the remaining symbols).
@@
    Satisfiability: verify if at least one model $S$ exists for a given theory $T$.
@@
    Logical entailment: verify whether a given statement $\phi$ is logically entailed by theory $T$.
@@
Detecting the correct inference to be executed is crucial to ensure that questions are correctly answered. To detect the intended inference from the a question, we fine-tuned a BERT sentence classifier [21] on a custom dataset, which contains over 1000 natural language questions and their corresponding logical reasoning task.
@@ 5.2 Information Extraction
Typically, in addition to a theory $T$, an interpretation is given for at least some of the symbols of $T$. For instance, the problem statement may be: “Given that Ann is 16 and Brit is 32, who is eligible for insurance?” Such information should be extracted and included as part of a structure on which the reasoning is performed (in this case, $age^{S}=\{\text{Ann}\mapsto 16,\text{Brit}\mapsto 32\}$). In addition, some other reasoning tasks also require additional information of similar complexity, e.g., optimization requires that we know which term $t$ should be minimized / maximized.
This information extraction task is made significantly easier by the fact that, at this point, we already know the vocabulary. We can therefore use language model grammars, such as llama.cpp GNBF grammar [11], to force the output to adhere to an automatically generated KB-specific grammar, ensuring that only correctly typed interpretations are generated. As we will show, this pruning of the output space allows us to achieves the same performance for this task with a small language model (SLM) as with a larger LLM. Since such a SLM can run on a standard laptop, this may significantly reduce computational costs, as well as enhance data security, by avoiding the transfer of potentially sensitive information to external servers.
@@
We currently do not attempt to automatically detect whether a specific question contains complex formulas or not, and thus requires an LLM, because we may wish to make this decision depending on the application. If we do not care about computational cost, we can just always invoke the LLM for optimal accuracy. Alternatively, we could base the decision whether to use an SLM or LLM purely on the detected reasoning task (e.g., Entailment is typically more likely to need an LLM), or we could even train a specific classifier to make the decision.
@@ 6.1 Validation
We first conduct a basic validation of VERUS-LM by checking (1) whether our use of a reasoning engine indeed leads to better perfomance on logical reasoning tasks than a stand-alone LLMs and (2) that the VERUS-LM pipeline is indeed able to correctly identify our eight different reasoning tasks from natural language text. Since existing datasets typically cover only one or at most a few reasoning tasks, we constructed our own dataset called DivLR in which all eight reasoning tasks are present. It consists of 115 questions covering six domains: investment Strategies, COVID restrictions, water Irrigation needs, a Handyman and two variants in the Body Mass Index (BMI) domain. The reason for having two BMI variants is that information about BMI (i.e., how to calculate it, the different risk levels associated to BMI values, etc.) probably occurs in LLM training data. In addition to our BMI domain which contains the well-known definition, we also defined a B* domain in which novel formulas and ranges for BMI were made up, similar to the experiment in [12].
We compare VERUS-LM to baseline language models on this dataset. We compare with both an SLM and an LLM, and consider two versions of our pipeline: one in which our information extraction is done by an SLM (V-SLM) and one in which this is done by an LLM (V-SLM). the LLM is Gemini 1.5 Pro [26] and the SLM is a quantised version of Phi 3.5 [2], a 3.8-billion-parameter language model whose inference cost is about 10,000 times lower than large models like GPT-4. All experiments were performed on a Mac M2 Pro CPU with 32GB of RAM.
Table 1: Results of VERUS-LM with SLM and LLM for information extraction and SLM and LLM separately as baseline. SLM = Phi 3.5 mini instruct, LLM = Gemini 1.5 Pro
@@
Table 1 shows that both V-SLM and V-LLM consistently outperform both the SLM and the LLM on their own. However, V-SLM performs significantly worse on the two BMI domains than on the other domain, and also significantly worse than V-LLM on these domains. A deeper analysis shows that the SLM typically refuses to simply extract the information that the reasoning engine needs, and instead tries to perform BMI-calculations itself. We conjecture that this is due to the BMI-related examples that are likely part of the data on which the SLM was trained.
[…]
@@
Table 2 confirms that our BERT classifier identifies the correct logical reasoning task with an average accuracy of 92.6%, performing consistently well on all tasks (with a small dip for Propagation). For the subsequent information extraction, we see that, for most of the reasoning tasks, V-LLM and S-SLM perform about equally well. The biggest difference is seen for Entailement. This is the only reasoning task in our DivLR dataset for which complex formulas need to be constructed at inference time, which V-SLM predictably struggles with (accuracy 29.6%). For the Propagation task, we also observe a difference in accuracy (-25.6%) between V-SLM and V-LLM. This is primarily due V-SLM struggling with the BMI domain, as already discussed.
Overall, we conclude that the proposed pipeline makes sense: (1) introducing a logical reasoning engine into the pipeline indeed leads to better performance than using only a language model; (2) different logical reasoning tasks can indeed by successfully identified by a simple BERT classifier; (3) all of the necessary information for the reasoning tasks can indeed be extracted with a small language model, as long as no complex formulas need to be constructed, and with the caveat that an SLM might be less robust when handling concepts that already occurred frequently in its training data.
@@ 6.2 Comparison to state-of-the-art
Table 3: Overview of datasets used for comparison with state-of-the-art
Dataset Example For inference # Options Reasoning #
PrOntoQA [22] “Stella is not hot” Data 2: $\top$, $\bot$ Satisfiability 500
ProofWriter [25] “The rabbit visits the cat” Data 3: $\top$$\bot,$ ? Propagation 600
FOLIO [13] “Marvin is neither a human nor from Mars” Formulas 3: $\top$, $\bot$, ? Logical entailment 204
LogicalDeduction [24] A)”Eve finished third.”, B) ”Max finished third”, $\ldots$ Only Data 3-5-7 Satisfiability 300
AR-LSAT [34] “What is the max #CDs on sale? A) two, B) six, …” Formulas 5 Multiple 231
Table 4: Comparison of the results of the neurosymbolic system. Results from original paper. Results from [32]. The results for Logic-LM are those with self refinements included, but without the LLM fallback in case of recurring syntax error.
System ProntoQA Proofwriter Folio
Logical
Deduction
AR-LSAT Average ST Dev
@@
We compare VERUS-LM against three LLM-based approaches and two neurosymbolic approaches for logical reasoning: (1) GPT-4 answering directly, (2) GPT-4 with Chain-of-Thought, (3) Linc [18] (only FOLIO and ProofWriter), (4) Logic-LM [19] and (5) SymbCot [32].
As shown in Table 4, all systems achieve a roughly similar performance on the first four benchmarks, with VERUS-LM scoring at most 5% worse than the best system on each, without relying on dataset-specific prompts. For the challenging AR-LSAT benchmark, VERUS-LM significantly outperforms all other methods, doing about 25% better than the second best system. We believe that this is due to (1) the generality of our approach, which does not use task-specific examples for in-context learning; (2) FO($\cdot$)’s expressiveness which allows for, e.g., straightforward representations of aggregates, such as “At least three CDs are on sale”: $\{c\in\textit{CD}:on\_sale(c)\}\geq 3$; (3) our support for different forms or reasoning, four of which were detected by the classifier and executed, as shown in Table 5.
Table 5: The distribution of reasoning tasks detected in AR-LSAT and VERUS-LM’s accuracy
Reasoning task Distribution (%) Exe_Acc (%)
@@ 6.3 Effect of Self-Refinement
No 90 95.7 86.2
ProofWriter Syntax 92.3 $\uparrow$ 2.3 95.7 == 88.3 $\uparrow$ 2.1
@@ 7 Limitations
Though VERUS-LM expands on the state of the art, it faces some limitations. Like the earlier papers in literature, VERUS-LM itself does not have a method for validating formula correctness. This is an inherent problem in this type of neuro-symbolic AI in general, as natural language can be ambiguous, and LLMs can always produce errors. To address this limitation, we are looking into bringing a domain expert in the loop without prior logic experience to validate the output [30].
Additionally, the expressiveness of FO($\cdot$) has limitations, making it unsuitable for, e.g., representing higher-order logic or for reasoning over OWA. Though VERUS-LM’s simulated OWA performed well on the datasets discussed in this work, it remains an approximation and thus will not be effective for all situations. Similarly, our pipeline’s semantic refinement step is also not always applicable, as there could be situations (such as reasoning over fake news) where we can expect unsatisfiability. Due to VERUS-LM’s relience on a logical reasoning engine, it also inherits some of its limitations, such as SAT problems in very large domains. Furthermore, though the KB-creation of VERUS-LM is based on a generic prompting pipeline, its accuracy in entirely different logical paradigms remains a uncertain.